A finance lease agreement calls for quarterly lease payments of $7,392 over a 10-year lease term, with the first payment on July 1, the beginning of the lease. The annual interest rate is 12%. Both the present value of the lease payments and the cost of the asset to the lessor are $176,000.

Required:
a. Prepare a partial amortization table up to the October 1 payment.
b. What would be the amount of interest expense (revenue) the lessee (lessor) would record in conjunction with the second quarterly payment on October 1?

A.

Date Lease Payment Effective Rate Decrease in balance Outstanding balance
July 1 ?
July 1 ? ? ?
October 1 ? ? ?

Part B

Interest expense (Lessee) ?
Interest Revenue (Lessor) ?
Answer #1

Answer:

a)Quarterly rate = 12 /4 = 3%     [4Quarters in a year]

During the period End
Date Carrying Value at beginning Payment Interest paid Principal repayment carrying value of lease at end
1July 176000 7392 0 7392 162000-6500=155500
1oct 168608 7392 168608*.03=5058 7392-5058= 2334 155500-1835= 153665

b)

Interest expense for lessee = $ 5058
Interest revenue for lessor = $ 5058